# Weak Memory Model Formalisms: Introduction and Survey

### v1

@@ 1. Introduction
To address this extra complication introduced by architectures and compiler optimisations, processor vendors and language designers define a “memory model”, a minimal set of constraints about the behaviour of interacting concurrent processes, typically with respect to the order of accesses of shared memory. If those constraints are too weak to enforce the logic of an algorithm, programmers may inject “fences” or related mechanisms that can restore the standard, sequential behaviour of a (segment of a) program. For instance, the use of release/acquire constraints is a well-known programming pattern for addressing weak memory model effects [127]. A release fence, ‘’, (in C, atomic_thread_fence(memory_order_release)) can be injected into the first process of (1), . In tandem with an acquire fence after the loop of the second process these ensure the intended value for will be read. However, excessive use of such mechanisms comes with performance penalties, which can be a significant issue for low-level systems code [250].
Since the early days of multiprocessors (a system consisting of multiple processors) to the present age of multicore processors (each processor consisting of multiple cores), the literature on weak memory models has developed in academia, the commercial sector, and standardisation committees. The issue affects hardware designers, compiler writers, application programmers, and language designers [17; 16]. As yet, no consensus on a particular formal representation has been settled upon, leaving programmers – in particular those requiring formal assurance for the code – with many choices. The company Arm has led the way in formalising their architecture and robustly investigating their design, including with publicly accessible tool support. Their formal model is validated on hardware through the use of thousands of small programs called “litmus tests”. However, the mechanisms for inferring properties of programs using their formalisation are complex, even for simple programs; as a result, many other representations have been considered.
@@
We aim to trace the historical development of processor design from the 1960s, through theoretical developments in the 1990s and up to the present day, which includes regular contributions in major publishing venues and refinements of commercial specifications. As a result, we were unable to find a methodical sequence of search terms of the type used in a recent survey on programming language memory models [219]. Hence, we do not claim full coverage, but intend that the range of works surveyed is indicative of the state of the field.
We do not specifically consider issues related to hardware or compiler construction: we assume that microarchitectural implementations of instruction set architectures faithfully match their specifications, and that compiler transformations correctly preserve their intended software memory model in the underlying architecture. Of course, this is not always the case, with problems being discovered in hardware [36; 38] and potential compiler mappings [284].
@@
As an introduction to formalisations and the range of representations that have been considered we present an operational (Section 4) representation based on write buffers (as defined in [261]), and and an axiomatic (Section 5) view that abstracts from architectural details and instead focusses on properties of traces of behaviours (based on work best exemplified in [36]). We use a representative fragment of Intel’s x86 as a running example of a memory model, and use the store buffering litmus test (Example 1.2) as the running example program. Because examples are traditionally the best way to explain weak memory effects, we keep the them separate from the main text, in highlighted boxes, to assist readers looking for instances over details.
@@ 1.1. Terminology and Conventions
Memory Models. The word “model” is heavily overloaded and potentially ambiguous.111 For example, a “memory model” can mean the structure of the memory address space. This usage, unrelated to concurrency, occurs in microarchitectures [133, p.72] as well as in C programming [189; 213]. In the setting of C, a memory model in this sense is sometimes called a “memory object model.” In this paper, a memory model222A memory model is also called memory consistency, consistency model and memory consistency model [220]. denotes the underlying set of possible behaviours of every concurrent program. Sequential Consistency is a memory model, and in fact the most simple one. When a program exhibits behaviour that is not compatible with Sequential Consistency, then the program is said to run under a weak (or relaxed) memory model.
@@ 1.2. Litmus Tests
Weak memory models are traditionally understood using small programs that highlight particular unexpected behaviours, typically called litmus tests. Given some initial state of the shared and local variables, one can “probe” a hardware or software memory model by checking whether a final state is reachable for a given litmus test. Examples 1.2–1.2 give several litmus tests, including the store buffering litmus test, which we use as a running example, and is one of the most common litmus tests referred to in the literature [209; 202]. Table 1 summarises the relationships between the key litmus tests and the common memory models.
@@ 2. The Development of Processors
| 1963 | The CDC 6600 [278] significantly improved processor speed by using a “scoreboard” to track which registers and locations are being accessed by instructions in the pipeline; instructions that have no dependencies may be parallelised. |
| 1967 | The first implementation of Tomasulo’s algorithm [279] was Model 91 in IBM’s System/360 family [37], which improved the utilisation of multiple arithmetic units. |
| 1980s | The Stanford DASH multiprocessor [188] brought together much existing research and influenced later processors. |
@@
| 2011 | RISC-V was released as an openly available specification [249], with a memory model based on Release Consistency (see Section 3.1). |
| 2017 | Based on research outcomes, as described in [36], the Arm specification explicitly became multicopy-atomic [240]. |
@@ 2.1. Instruction-Level Parallelism
The execution of a single instruction may involve multiple micro-steps, including to access registers, calculate values, or read and write main memory. These latter micro-steps, which we call “memory operations”, can include complex interactions with the cache system, or across PCI devices. Any one of these or many other aspects of computer hardware [147; 230; 212] could potentially influence the actual order of execution of the original sequence of compiled instructions.
Focussing at the local level, the performance of a processor can be improved in part by executing independent micro-steps in parallel. Techniques in this vein are collectively termed instruction-level parallelism, and have been developed since the early days of computing [246]. Straightforward cases include parallelising loads of distinct locations in memory, and parallelising independent calculations to utilise multiple arithmetic and logic units (ALUs) on a single chip. The decision whether or not to parallelise is guided by the observable effect on the sequential semantics – if the programmer could potentially observe the parallelism on a single chip, then it should not be allowed (see Example 2.1). Early examples include the “scoreboarding” technique of the CDC 6600 computer [278] and Tomasulo’s algorithm [279]. Parallelism can also be employed when accessing distinct locations in memory (see Example 2.1).
[t]
@@
[t]
These “reordering” principles were extended by Intel’s Pentium processors in the 1990s to include speculative execution: while waiting for a branch condition to be resolved to true or false (e.g., while waiting for a value to return from memory), a processor may predict the result and begin transiently executing down the corresponding path. This transient execution may include executing memory operations associated with loads (but not stores) out-of-order with instructions before the branch point. If the prediction is later seen to be correct then the effects of the transient execution are committed; otherwise, the transient calculations are discarded and fetching restarts at the alternative (not-predicted) path. Where processor speeds are much faster than memory latency, speculative execution provides a substantial performance gain. Such optimisations have led to significant security vulnerabilities, exposed by attacks such as Spectre [171] and Meltdown [193], which were only discovered decades after speculative execution and other optimisation mechanisms had become standard.
@@ 2.2. Fences and Memory Ordering Constraints
Many different types of fences exist; a full fence ensures that all memory operations before it have been issued, and prevents all later memory operations after it from executing early (see Example 2.2: x86 provides mfence, Arm provides dmb, and RISC-V provides fence.rw.rw). Some processors come with more fine-grained fences; such a specialised fence may only affect specific types of memory operations (loads or stores), or may only affect instructions before or after it. Memory ordering constraints may also be used to enforce ordering between otherwise independent instructions; the classic examples are the release and acquire constraints pioneered by Release Consistency [127] (see Example 2.2 and Section 3.1: as an example, Arm provides both LDAR and LDAPR instructions for acquire loads).
[t]
@@ 2.4. Multicopy Atomicity
[ht]
Currently, IBM’s Power series is the only commercially available CPU architecture that lacks multicopy atomicity (the NVIDIA PTX GPU also lacks multicopy atomicity [224]). This is due to its cache coherence system, which effectively allows individual cores to communicate point-to-point via shared caches, before writes are flushed to the shared memory.
The lack of multicopy atomicity creates an extra level of complexity for describing memory models, as well as understanding their effects on code – both at the assembler-level as well as in high-level languages to be compiled to the hardware. To this end, Arm officially stipulates multicopy atomicity for their architectures from Version 8 in 2017 [240]. (In fact, Arm processors effectively have always been multicopy-atomic. Although multicopy atomicity was not explicitly stipulated by earlier versions of Arm, no manufacturer of Arm processors ever produced an implementation that lacked multicopy atomicity [240].) Similarly, RISC-V insists on multicopy atomicity, and Intel’s x86 architectures are also multicopy-atomic.
In practice, the complexity of dealing with an architecture that lacks multicopy atomicity leads programmers to insert enough fences to restore multicopy atomicity (or indeed Sequential Consistency). We therefore do not consider the lack of multicopy atomicity in this paper; a description of the behaviour of Power is given in [122], and the difficulties of verification under Power’s memory model are discussed in [94].
@@ 3. The Development of Memory Models
| 1990s | [263] | [92] | |||||
| 2000s | [227] | [208] | [55] | ||||
| 2010s | [122] | [249] | [50] | ||||
| 2020s | [27] |
@@ 3.1. Early Memory Models
Lamport [184] made the first exploration on the notion of the “intuitive correctness” of concurrent programs in the 1970s, and formalised the concept of what later came to be known as Sequential Consistency. From the late 1980s onwards there were considerable efforts in defining the semantics of various instruction-level parallelism techniques, such as pipelining and out-of-order execution (see Section 2) [65; 145; 160; 262]. This saw the proposal of many theoretical memory models, aiming to categorise and unify observed and potential behaviours. Examples include Weak Consistency [116], PRAM Consistency [194], Release Consistency [127], Cache Consistency [134], Processor Consistency [134], and Causal Consistency [21]. These memory models were unified in 2004 by Steinke and Nutt [272], who used a mathematical formalism to compare these memory models and to arrange them in a lattice. (In 2018, these theoretical models were again unified and compared [256; 257].)
@@ 3.2. Hardware Memory Models
The Total Store Order (TSO) and Partial Store Order (PSO) memory models were defined for the SPARC architecture of Sun Microsystems [263]. As the name suggests, TSO keeps the writes of each thread in program order, but a read that accesses some location can complete before an earlier write of the same thread to a location different from . PSO relaxes this further to maintain order only on writes to the same variable. SPARC also defines a Relaxed Memory Order (RMO) model, which “places no ordering constraints on memory references beyond those required for processor self-consistency” [271]. The TSO model of SPARC is formalised in Isabelle/HOL [150].
@@
The x86 family of architectures is widely implemented by various companies such as Intel and AMD. Aiming to devise a rigorous specification of the behaviour of x86, Sarkar et al. [255] developed the memory model x86-CC based on Causal Consistency (see Section 3.1). However, x86-CC was later found not to match the actual behaviour of the hardware in all cases. Consequently, Owens et al. [227] developed x86-TSO, after observing similarities between the behaviour of x86 and the TSO memory model of SPARC. To this date, x86-TSO is widely accepted as the correct description of the behaviour of x86 architectures. In [227], x86-TSO was introduced in terms of two formulations, one operational and one axiomatic. The operational formulation uses a write buffer to model the weak memory effects, and is influential on many later formalisations of memory models; we provide a version of it in Section 4. The two key papers [227; 255] that developed x86-TSO are summarised by Sewell et al. [260].
@@
The Power family of processors by IBM has a very relaxed memory model that lacks multicopy atomicity (see Section 2.4), permitting many possible behaviours. An axiomatic formulation for Power’s memory model was first defined by Corella et al. [92], which was improved by Adir et al. [15] a decade later based on observed behaviours of litmus tests. Alglave et al. [33, 35, 36] also studied the Power memory model based on an axiomatic definition. An alternative, operational approach to formalisation of Power was given by Sarkar et al. [254].
@@
Whereas the desktop and laptop markets are dominated by x86, as of 2025 the mobile device market is dominated by Arm architectures. Arm architectures previously did not forbid non-multicopy-atomic behaviours, but after a major revision in 2017, Arm has mandated multicopy atomicity from Version 8 onwards [240; 27]. There have been numerous works on specifying versions of the Arm memory model using different formalisms, such as [122]. The development history is included in the latest official paper on the Arm memory model [27].
@@
The RISC-V architecture [249], first released in 2010, is an open-source architecture that attempts to remove some of the problems associated with commercial-in-confidence manufacturing. Its memory model is based on experiences at Arm, and includes many recent developments in weak memory theory, including support for release-acquire primitives.
@@
Recent research effort has focussed on GPUs [24; 270; 198] and FPGAs [151], and their combination with CPUs [130]. GPUs in particular focus on significant parallelisation of processing power, and hence would benefit from formalisation; NVIDIA has provided a rigorous description of its memory model [224], which lacks multicopy atomicity (see Section 2.4). Earlier formalisations determined a type of total-store-order model [24], while more recent formalisations suggest a weaker model again [198].
@@
Intel’s Itanium processors [14] never became as commercially popular as other processors, and have since ceased production. Similarly, DEC’s Alpha processors were prominent for a time, but have also ceased production.

### v2

@@ 1. Introduction
To address this extra complication introduced by architectures and compiler optimisations, processor vendors and language designers define a “memory model”, a minimal set of constraints about the behaviour of interacting concurrent processes, typically with respect to the order of accesses of shared memory. If those constraints are too weak to enforce the logic of an algorithm, programmers may inject “fences” or related mechanisms that can restore the standard, sequential behaviour of a (segment of a) program. For instance, the use of release/acquire constraints is a well-known programming pattern for addressing weak memory model effects [Gharachorloo et al., 1990]. A release fence, ‘’, (in C, atomic_thread_fence(memory_order_release)) can be injected into the first process of (1), . In tandem with an acquire fence after the loop of the second process these ensure the intended value for will be read. However, excessive use of such mechanisms comes with performance penalties, which can be a significant issue for low-level systems code [Ritson and Owens, 2016].
Since the early days of multiprocessors (a system consisting of multiple processors) to the present age of multicore processors (each processor consisting of multiple cores), the literature on weak memory models has developed in academia, the commercial sector, and standardisation committees. The issue affects hardware designers, compiler writers, application programmers, and language designers [Adve and Gharachorloo, 1996; Adve and Boehm, 2010]. As yet, no consensus on a particular formal representation has been settled upon, leaving programmers – in particular those requiring formal assurance for the code – with many choices. The company Arm has led the way in formalising their architecture and robustly investigating their design, including with publicly accessible tool support. Their formal model is validated on hardware through the use of thousands of small programs called “litmus tests”. However, the mechanisms for inferring properties of programs using their formalisation are complex, even for simple programs; as a result, many other representations have been considered.
@@
We aim to trace the historical development of processor design from the 1960s, through theoretical developments in the 1990s and up to the present day, which includes regular contributions in major publishing venues and refinements of commercial specifications. As a result, we were unable to find a methodical sequence of search terms of the type used in a recent survey on programming language memory models [Moiseenko et al., 2021]. Hence, we do not claim full coverage, but intend that the range of works surveyed is indicative of the state of the field.
We do not specifically consider issues related to hardware or compiler construction: we assume that microarchitectural implementations of instruction set architectures faithfully match their specifications, and that compiler transformations correctly preserve their intended software memory model in the underlying architecture. Of course, this is not always the case, with problems being discovered in hardware [Alglave et al., 2014b; Arm Ltd., 2011] and potential compiler mappings [Trippel et al., 2018].
@@
As an introduction to formalisations and the range of representations that have been considered we present an operational (Section 4) representation based on write buffers (as defined in [Sewell et al., 2010b]), and and an axiomatic (Section 5) view that abstracts from architectural details and instead focusses on properties of traces of behaviours (based on work best exemplified in [Alglave et al., 2014b]). We use a representative fragment of Intel’s x86 as a running example of a memory model, and use the store buffering litmus test (Example 1.2) as the running example program. Because examples are traditionally the best way to explain weak memory effects, we keep the them separate from the main text, in highlighted boxes, to assist readers looking for instances over details.
@@ 1.1. Terminology and Conventions
Memory Models. The word “model” is heavily overloaded and potentially ambiguous.111For example, a “memory model” can mean the structure of the memory address space. This usage, unrelated to concurrency, occurs in microarchitectures [González et al., 2011, p.72] as well as in C programming [Lepigre et al., 2022; Memarian et al., 2016]. In the setting of C, a memory model in this sense is sometimes called a “memory object model.” In this paper, a memory model222A memory model is also called memory consistency, consistency model and memory consistency model [Nagarajan et al., 2020]. denotes the underlying set of possible behaviours of every concurrent program. Sequential Consistency is a memory model, and in fact the most simple one. When a program exhibits behaviour that is not compatible with Sequential Consistency, then the program is said to run under a weak (or relaxed) memory model.
@@ 1.2. Litmus Tests
Weak memory models are traditionally understood using small programs that highlight particular unexpected behaviours, typically called litmus tests. Given some initial state of the shared and local variables, one can “probe” a hardware or software memory model by checking whether a final state is reachable for a given litmus test. Examples 1.2–1.2 give several litmus tests, including the store buffering litmus test, which we use as a running example, and is one of the most common litmus tests referred to in the literature [Maranget et al., 2012; Mador-Haim et al., 2011]. Table 1 summarises the relationships between the key litmus tests and the common memory models.
@@ 2. The Development of Processors
| 1963 | The CDC 6600 [Thornton, 1964] significantly improved processor speed by using a “scoreboard” to track which registers and locations are being accessed by instructions in the pipeline; instructions that have no dependencies may be parallelised. |
| 1967 | The first implementation of Tomasulo’s algorithm [Tomasulo, 1967] was Model 91 in IBM’s System/360 family [Anderson et al., 1967], which improved the utilisation of multiple arithmetic units. |
| 1980s | The Stanford DASH multiprocessor [Lenoski et al., 1992] brought together much existing research and influenced later processors. |
@@
| 2011 | RISC-V was released as an openly available specification [RISC-V Foundation, 2019], with a memory model based on Release Consistency (see Section 3.1). |
| 2017 | Based on research outcomes, as described in [Alglave et al., 2014b], the Arm specification explicitly became multicopy-atomic [Pulte et al., 2017]. |
@@ 2.1. Instruction-Level Parallelism
The execution of a single instruction may involve multiple micro-steps, including to access registers, calculate values, or read and write main memory. These latter micro-steps, which we call “memory operations”, can include complex interactions with the cache system, or across PCI devices. Any one of these or many other aspects of computer hardware [Hennessy and Patterson, 2011; Patterson and Hennessy, 2016; McKenney, 2010] could potentially influence the actual order of execution of the original sequence of compiled instructions.
Focussing at the local level, the performance of a processor can be improved in part by executing independent micro-steps in parallel. Techniques in this vein are collectively termed instruction-level parallelism, and have been developed since the early days of computing [Rau and Fisher, 1993]. Straightforward cases include parallelising loads of distinct locations in memory, and parallelising independent calculations to utilise multiple arithmetic and logic units (ALUs) on a single chip. The decision whether or not to parallelise is guided by the observable effect on the sequential semantics – if the programmer could potentially observe the parallelism on a single chip, then it should not be allowed (see Example 2.1). Early examples include the “scoreboarding” technique of the CDC 6600 computer [Thornton, 1964] and Tomasulo’s algorithm [Tomasulo, 1967]. Parallelism can also be employed when accessing distinct locations in memory (see Example 2.1).
@@
These “reordering” principles were extended by Intel’s Pentium processors in the 1990s to include speculative execution: while waiting for a branch condition to be resolved to true or false (e.g., while waiting for a value to return from memory), a processor may predict the result and begin transiently executing down the corresponding path. This transient execution may include executing memory operations associated with loads (but not stores) out-of-order with instructions before the branch point. If the prediction is later seen to be correct then the effects of the transient execution are committed; otherwise, the transient calculations are discarded and fetching restarts at the alternative (not-predicted) path. Where processor speeds are much faster than memory latency, speculative execution provides a substantial performance gain. Such optimisations have led to significant security vulnerabilities, exposed by attacks such as Spectre [Kocher et al., 2019] and Meltdown [Lipp et al., 2018], which were only discovered decades after speculative execution and other optimisation mechanisms had become standard.
@@ 2.2. Fences and Memory Ordering Constraints
Many different types of fences exist; a full fence ensures that all memory operations before it have been issued, and prevents all later memory operations after it from executing early (see Example 2.2: x86 provides mfence, Arm provides dmb, and RISC-V provides fence.rw.rw). Some processors come with more fine-grained fences; such a specialised fence may only affect specific types of memory operations (loads or stores), or may only affect instructions before or after it. Memory ordering constraints may also be used to enforce ordering between otherwise independent instructions; the classic examples are the release and acquire constraints pioneered by Release Consistency [Gharachorloo et al., 1990] (see Example 2.2 and Section 3.1: as an example, Arm provides both LDAR and LDAPR instructions for acquire loads).
@@ 2.4. Multicopy Atomicity
Currently, IBM’s Power series is the only commercially available CPU architecture that lacks multicopy atomicity (the NVIDIA PTX GPU also lacks multicopy atomicity [NVIDIA Ltd, 2024]). This is due to its cache coherence system, which effectively allows individual cores to communicate point-to-point via shared caches, before writes are flushed to the shared memory.
The lack of multicopy atomicity creates an extra level of complexity for describing memory models, as well as understanding their effects on code – both at the assembler-level as well as in high-level languages to be compiled to the hardware. To this end, Arm officially stipulates multicopy atomicity for their architectures from Version 8 in 2017 [Pulte et al., 2017]. (In fact, Arm processors effectively have always been multicopy-atomic. Although multicopy atomicity was not explicitly stipulated by earlier versions of Arm, no manufacturer of Arm processors ever produced an implementation that lacked multicopy atomicity [Pulte et al., 2017].) Similarly, RISC-V insists on multicopy atomicity, and Intel’s x86 architectures are also multicopy-atomic.
In practice, the complexity of dealing with an architecture that lacks multicopy atomicity leads programmers to insert enough fences to restore multicopy atomicity (or indeed Sequential Consistency). We therefore do not consider the lack of multicopy atomicity in this paper; a description of the behaviour of Power is given in [Flur et al., 2016], and the difficulties of verification under Power’s memory model are discussed in [Coughlin et al., 2022].
@@ 3. The Development of Memory Models
| 1990s | [Sindhu et al., 1992] | [Corella et al., 1993] | |||||
| 2000s | [Owens et al., 2009] | [Manson et al., 2005] | [Boehm, 2007] | ||||
| 2010s | [Flur et al., 2016] | [RISC-V Foundation, 2019] | [Batty et al., 2011] | ||||
| 2020s | [Alglave et al., 2021] |
@@ 3.1. Early Memory Models
Lamport [1979] made the first exploration on the notion of the “intuitive correctness” of concurrent programs in the 1970s, and formalised the concept of what later came to be known as Sequential Consistency. From the late 1980s onwards there were considerable efforts in defining the semantics of various instruction-level parallelism techniques, such as pipelining and out-of-order execution (see Section 2) [Burch and Dill, 1994; Harman, 2007; Jones et al., 2002; Shasha and Snir, 1988]. This saw the proposal of many theoretical memory models, aiming to categorise and unify observed and potential behaviours. Examples include Weak Consistency [Dubois et al., 1986], PRAM Consistency [Lipton and Sandberg, 1988], Release Consistency [Gharachorloo et al., 1990], Cache Consistency [Goodman, 1991], Processor Consistency [Goodman, 1991], and Causal Consistency [Ahamad et al., 1995]. These memory models were unified in 2004 by Steinke and Nutt [2004], who used a mathematical formalism to compare these memory models and to arrange them in a lattice. (In 2018, these theoretical models were again unified and compared [Senftleben and Schneider, 2018a, b].)
@@ 3.2. Hardware Memory Models
The Total Store Order (TSO) and Partial Store Order (PSO) memory models were defined for the SPARC architecture of Sun Microsystems [Sindhu et al., 1992]. As the name suggests, TSO keeps the writes of each thread in program order, but a read that accesses some location can complete before an earlier write of the same thread to a location different from . PSO relaxes this further to maintain order only on writes to the same variable. SPARC also defines a Relaxed Memory Order (RMO) model, which “places no ordering constraints on memory references beyond those required for processor self-consistency” [SPARC International Inc., 1994]. The TSO model of SPARC is formalised in Isabelle/HOL [Hóu et al., 2021].
@@
The x86 family of architectures is widely implemented by various companies such as Intel and AMD. Aiming to devise a rigorous specification of the behaviour of x86, Sarkar et al. [2009] developed the memory model x86-CC based on Causal Consistency (see Section 3.1). However, x86-CC was later found not to match the actual behaviour of the hardware in all cases. Consequently, Owens et al. [2009] developed x86-TSO, after observing similarities between the behaviour of x86 and the TSO memory model of SPARC. To this date, x86-TSO is widely accepted as the correct description of the behaviour of x86 architectures. In [Owens et al., 2009], x86-TSO was introduced in terms of two formulations, one operational and one axiomatic. The operational formulation uses a write buffer to model the weak memory effects, and is influential on many later formalisations of memory models; we provide a version of it in Section 4. The two key papers [Owens et al., 2009; Sarkar et al., 2009] that developed x86-TSO are summarised by Sewell et al. [2010a].
@@
The Power family of processors by IBM has a very relaxed memory model that lacks multicopy atomicity (see Section 2.4), permitting many possible behaviours. An axiomatic formulation for Power’s memory model was first defined by Corella et al. [1993], which was improved by Adir et al. [2003] a decade later based on observed behaviours of litmus tests. Alglave et al. [2010, 2012, 2014b] also studied the Power memory model based on an axiomatic definition. An alternative, operational approach to formalisation of Power was given by Sarkar et al. [2011].
@@
Whereas the desktop and laptop markets are dominated by x86, as of 2025 the mobile device market is dominated by Arm architectures. Arm architectures previously did not forbid non-multicopy-atomic behaviours, but after a major revision in 2017, Arm has mandated multicopy atomicity from Version 8 onwards [Pulte et al., 2017; Alglave et al., 2021]. There have been numerous works on specifying versions of the Arm memory model using different formalisms, such as [Flur et al., 2016]. The development history is included in the latest official paper on the Arm memory model [Alglave et al., 2021].
@@
The RISC-V architecture [RISC-V Foundation, 2019], first released in 2010, is an open-source architecture that attempts to remove some of the problems associated with commercial-in-confidence manufacturing. Its memory model is based on experiences at Arm, and includes many recent developments in weak memory theory, including support for release-acquire primitives.
@@
Recent research effort has focussed on GPUs [Alglave et al., 2015; Sorensen and Donaldson, 2016; Lustig et al., 2019] and FPGAs [Iorga et al., 2021], and their combination with CPUs [Goens et al., 2023]. GPUs in particular focus on significant parallelisation of processing power, and hence would benefit from formalisation; NVIDIA has provided a rigorous description of its memory model [NVIDIA Ltd, 2024], which lacks multicopy atomicity (see Section 2.4). Earlier formalisations determined a type of total-store-order model [Alglave et al., 2015], while more recent formalisations suggest a weaker model again [Lustig et al., 2019].
@@
Intel’s Itanium processors [Abe and Maeda, 2017] never became as commercially popular as other processors, and have since ceased production. Similarly, DEC’s Alpha processors were prominent for a time, but have also ceased production.